OpenStudy (finn):

Solve the system of linear equations by substitution y = 5x + 6 y = -2x - 8

OpenStudy (anonymous):

cant explain it but i can give you an example: equation 1: 14x+2y=20 equation 2: 5x-y=8 First step is to get on of the varibles to cancel out so leave the first eqation alone put change the second in this problem so: 5x-y=8 becomes 2(5x-y=8) which equals 10x-2y=16 the next step is to subtract them like this: 14x+2y=20 10x-2y=16 the +2y and the -2y cancel out so the answer then would be 4x=4 which means 1=x then you plug the answer 1 in for x into on of the equations to find y so: 5(1)-y=8 = 5-y=8 =-y=3 =y=-3 so the system answer is (1,-3)

OpenStudy (anonymous):

i just gave you an example

OpenStudy (anonymous):

-5x-6 = 5 Add 5 6 to both sides. -5x = 11 Divide both sides by -5 x = -11/5 Now substitute this for x into either of the two original equations and solve for y. Let's use the second equation: y-5x = 3 Substitute x = -11/5 y-5(-11/5) = 3 Simplify. y+11 = 3 Subtract 11 from both sides. y = - 8 The solution is: (-11/5, -8) Let's check the graph of the two lines represented by the two equations. The solution will be seen as the point of intersection of the two lines.

OpenStudy (anonymous):

Which set of data has 3 modes? 33, 34, 31, 34, 39, 35, 37, 31, 39, 30 31, 34, 31, 34, 39, 35, 37, 31, 34, 35 32, 35, 31, 36, 41, 40, 37, 30, 42, 29 32, 35, 36, 35, 32, 36, 32, 36, 35, 32